OpenStudy (anonymous):

how about f(x)=(x^3 -8)/(x-2) when x approach 2

OpenStudy (anonymous):

should i factorise first?

OpenStudy (anonymous):

Yes!

zepdrix (zepdrix):

Remember how to factor out a term when you have the sum or difference of cubes? :)

OpenStudy (anonymous):

its factorise of x^3-8 = (x-2)(x^2-1)

OpenStudy (anonymous):

i can't factorise it..

zepdrix (zepdrix):

8 = 2^3 Difference of cubes looks like this: \[a^3-b^3=(a-b)(a^2+ab+b^2)\] I think that's how it goes at least XD I'll check to make sure..

zepdrix (zepdrix):

Ya :D So you have x^3 - 2^3, a=x b=2 Can you factor it using the formula i posted? c:

zepdrix (zepdrix):

Confused? :o

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

i got d answer..thanks